AnimeGhoul8863:

Rachel can make 3 bracelets in an hour. Oliver can make only 2 bracelets in an hour, but he already has 5 completed bracelets. Explain to Rachel how she can use a system of equations to determine when she will have the same number of bracelets as Oliver. Use complete sentences.

AnimeGhoul8863:

@dude e.e could yew help meh

Bearclaws72:

Since rachel makes 3 bracelets an hour, her pinear equstion should be y=3h. 3 represents the rate of change and h represents the number of hours as a variable. Oliver's linear equation is y=2h+5 because 2 is the rate of chabge and 5 is the b value or the y intercept in a linear equation y=mx+b. To find when both have the same number of bracelets, we find the point of intersection... 2h+5=3h 5=3h-2h 5=h Therefore in 5 hours, both will have the same amount of bracelets.

AnimeGhoul8863:

Rachel can use the system of equations to help her get the same amount of bracelets as Oliver by turning Rachel's 3 bracelets into y=3h This will represent as rate of change (3) and Hours it takes to complete (h). This is the first equation out of 2. Oliver's Equation is y=2h+5 meaning the same this as the first equation but this time adding 5 as our b value because he already has 5 bracelets already made. Now in order to figure out how long it takes to make as much bracelets as oliver you need to find the intersection now putting it all together it should look like this 2h + 5 = 3h 5 = 3h - 2h 5 = h

AnimeGhoul8863:

^ is this right ?

Bearclaws72:

Yes

AnimeGhoul8863:

thank yew e.e

Bearclaws72:

Take back the medal

AnimeGhoul8863:

Biss just cause u call meh a loser im not taking it back -.-

AnimeGhoul8863:

u da loser

AnimeGhoul8863:

u know what i take it back u dun deserve